Please Help!
Compute the specified quantity.
$4,000 loan, taken now, with a simple interest rate of 6% per year, will require a total repayment of $5,440. At what time t will the loan mature?
Thank You

Answer by stanbon(75887)

Compute the specified quantity. $4,000 loan, taken now, with a simple interest rate of 6% per year, will require a total repayment of $5,440.
At what time t will the loan mature?
---------------------
A(t) = P(1+rt)
5440 = 4000(1+0.06t)
-------
1 + 0.06t = 1.36
0.06t =0.36
----
t = 0.36/0.06
t = 6 years
---------------
Cheers,
Stan H.
-------------